# Extracting Interpretable Logic Rules from Graph Neural Networks

### v1

@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as powerful tools for modeling and analyzing graph-structured data, achieving remarkable performance across diverse domains, including drug discovery (Xiong et al., 2021; Liu et al., 2022; Sun et al., 2020), fraud detection (Rao et al., 2021), and recommender systems (Chen et al., 2022). Despite their success, GNNs share the black-box nature inherent to the neural network family, which poses challenges to their further development in high-reliability applications such as healthcare (Amann et al., 2020; Bussmann et al., 2021).
@@
To this end, several explainability methods have been developed to uncover the inner decision-making mechanisms of GNNs. However, most of these methods are limited to providing local explanations tailored to specific input instances or rely on interpretations based on input feature attributions (Pope et al., 2019; Ying et al., 2019; Vu & Thai, 2020; Lucic et al., 2022; Tan et al., 2022). Another line of research focuses on global explanations that describe the overall behavior of models (Xuanyuan et al., 2023; Azzolin et al., 2023). These approaches offer more human-readable and precise explanations by leveraging logical formulas and interpretable concepts. However, they have limitations, such as relying on predefined concepts and generating rules that explain only a limited set of patterns within each class without effectively distinguishing between these classes.
@@
To this end, we propose LogicXGNN, a novel framework that extracts interpretable logic rules to explain the internal reasoning process of GNNs while maintaining classification accuracy comparable to the original model. LogicXGNN ensures interpretability by grounding the hidden predicates of the rules into the input space using decision trees, which are both computationally efficient and data-driven. To the best of our knowledge, LogicXGNN is the first interpretable, rule-based functional equivalent of GNNs. Additionally, it can function as a generative model, with notable potential in fields such as drug design and knowledge discovery. We demonstrate the effectiveness of our approach through experiments on real-world datasets, including the MUTAG and BBBP datasets. In summary, we make the following contributions:
  • We propose LogicXGNN, a novel post-hoc approach for extracting interpretable logic rules from graph neural networks. Our method incurs minimal computational overhead, comparable to training decision trees, and is model-agnostic, compatible with a wide range of GNN architectures.
  • LogicXGNN not only generates discriminative rules that serve as a rule-based counterpart to the original model, achieving comparable classification accuracy, but also facilitates knowledge discovery.
@@
  • As LogicXGNN makes decision-making in GNNs transparent with descriptive rules, it can be used as a generative model for creating graph instances, holding significant potential in fields such as drug design.
@@ 2.1 Graph Neural Networks
$\mathbf{h}_{v}^{l+1}=\text{UPD}\left(\mathbf{h}_{v}^{l},\text{AGG}\left(\left\{\text{MSG}(\mathbf{h}_{v}^{l},\mathbf{h}_{u}^{l})\mid\mathbf{A}_{uv}=1\right\}\right)\right),$
where $\mathbf{h}_{v}^{0}=\mathbf{X}_{v}$ is the initial feature vector of node $v$, and $\mathbf{h}_{v}^{l}$ represents the final node embedding at the final layer $L$. The update function UPD, aggregation operation AGG, and message function MSG define the architecture of a GNN. For instance, Graph Convolutional Networks (GCN) (Kipf & Welling, 2017) use an identity message function, mean aggregation, and a weighted update. Non-linearities are incorporated into the model through activation functions, typically the Rectified Linear Unit (ReLU), applied as: $\mathbf{h}_{v}^{l}=\text{ReLU}(\mathbf{h}_{v}^{l})$. In essence, the GNN $\mathcal{M}$ aggregates information from both the feature space and the topological structure of $G$ to compute deep hidden embeddings for nodes. These embeddings are then optimized for specific downstream tasks, such as node or graph classification.
Node Classification
We assume a label function $f:\mathcal{V}\rightarrow\{1,\dots,C\}$ that assigns one of $C$ classes to each node in $\mathcal{V}$. To approximate $f$, we define a GNN model $\mathcal{M}$ for node classification by passing the final node embeddings $\mathbf{h}_{v}^{(L)}$ through a softmax function for each node $v$.
@@
Similarly, suppose we have a set of graphs $\mathcal{G}$ and a label function $f:\mathcal{G}\rightarrow\{1,\dots,C\}$ that assigns one of $C$ classes to each graph in $\mathcal{G}$. To approximate $f$, we define a GNN model $\mathcal{M}$ for graph classification by passing the graph embeddings $\mathbf{h}_{G}^{L}$ to a fully connected layer followed by a softmax function. Here, the graph embeddings are commonly computed by taking the mean of all node embeddings in the graph $\mathbf{h}_{G}^{L}:=\text{mean}(\mathbf{h}_{v}^{L}\mid v\in V_{G})$ through the operation global_mean_pooling.
@@ 2.2 First-Order Logic Rules for GNN Interpretability
First-order logic (FOL), also known as predicate logic, is highly interpretable to humans, making it an excellent tool for explaining the behaviour of neural networks (Zhang et al., 2021). In this paper, our proposed framework, LogicXGNN, aims to elucidate the inner decision-making process of a GNN $\mathcal{M}$ using a Disjunctive Normal Form (DNF) formula $\phi_{M}$. The formula $\phi_{M}$ is a logical expression that can be described as a disjunction of conjunctions (OR of ANDs) over a set of predicates $P$, where each $p_{j}$ represents a logical condition or property defined on the graph structure $\mathbf{A}$ and input features $\mathbf{X}$. Importantly, $\phi_{M}$ incorporates the universal quantifier ($\forall$), providing a global explanation that is specific to a class of instances. This makes $\phi_{M}$ a rule-based model that is functionally equivalent to $\mathcal{M}$ a significant advantage not typically offered by other explanation work, such as attribute-based approaches. To provide a concrete example of $\phi_{M}$, suppose the model $M$ is trained to determine whether a molecule $G$ is soluble. In this case, we may extract the following logical rules $\phi_{M}$ from $M$:
$\forall G,\,(p_{1}(G)\land p_{2}(G))\lor(p_{3}(G)\land p_{4}(G))\Rightarrow\text{label}(G,s),$
where $p_{1}(G)$ represents “The molecule contains a hydroxyl group (-OH)”, $p_{2}(G)$ represents “The molecule has a ring structure”, $p_{3}(G)$ represents “The molecule contains a carbonyl group (C=O)”, $p_{4}(G)$ represents “The molecule has a high degree of branching”, and $\text{label}(G,s)$ is a predicate indicating that graph $G$ is assigned the class label $s$ (“soluble”). Similarly, for node classification tasks, predicates $\{p_{j}(v)\}$ encode logical conditions and properties on nodes’ neighbourhoods and features.
@@
We show how LogicXGNN addresses these challenges in Section 3. Our key observation is that neural networks, including GNNs, exhibit class-specific activation patterns in their hidden representational spaces, capturing diverse instances within the same class (Yosinski et al., 2015; Gopinath et al., 2019; Geng et al., 2023, 2024). The central idea of LogicXGNN is to leverage these hidden patterns as predicates and then identify logical structures in a data-driven manner, ultimately grounding the predicates into the input space. This design choice aligns with the spirit of a recent work, NeuroLogic (Geng et al., 2025), which shows promising results in explaining fully connected networks and convolutional neural networks. Interested readers can refer to the original paper for further details.
@@ 3 The LogicXGNN Framework
This section provides a detailed discussion of our methodology and design choices in LogicXGNN. For ease of discussion, we divide LogicXGNN into three critical sub-problems. Figure LABEL:fig:overview illustrates an overview of the LogicXGNN framework.
@@ 3.1 Identifying Hidden Predicates $P$ for $\phi_{M}$
We start by discussing the identification of hidden predicates for graph classification tasks. In fact, we define hidden predicates at the node level, so it naturally works for node classification tasks, as will be clearly explained later.
As previously mentioned, the desired predicates $P$ should capture commonly shared patterns in both graph structures $\mathbf{A}$ and hidden embeddings $\mathbf{h}^{L}$ across a set of instances in the context of GNNs. While graph structure information can be encoded into hidden embeddings, it often becomes indistinguishable due to oversmoothing during the message-passing process. To mitigate this, we explicitly model common patterns in graph structures.
@@
$\text{Pattern}_{\text{struct}}(v)=\text{Hash}\big{(}\text{ReceptiveField}(v,\mathbf{A},L)\big{)}.$
Next, we discuss common patterns in the hidden embeddings. During the training of GNNs for classification tasks, the hidden embeddings are optimized to differentiate between classes. Empirically, we find that a small subset of specific dimensions in the final-layer embeddings $\mathbf{h}_{G}^{L}$ is sufficient to distinguish instances from different classes when using appropriate thresholds, often achieving similar accuracy to the original neural networks. Similar observations have been reported in (Geng et al., 2025).
@@
$\text{DecisionTree}(\{\mathbf{h}_{G}^{L}\mid G\in\mathcal{G}\},\mathbf{Y})\rightarrow(K,T)$
@@
$\mathcal{I}_{k}({\mathbf{h}_{v}^{L}})=\begin{cases}1&\text{if }\mathbf{h}_{v}^{L}[k]\geq T_{k}\\ 0&\text{if }\mathbf{h}_{v}^{L}[k]<T_{k}\end{cases}$
@@
$\text{Pattern}_{\text{emb}}(v)=\left[\mathcal{I}_{1}(\mathbf{h}_{v}^{L}),\mathcal{I}_{2}(\mathbf{h}_{v}^{L}),\dots,\mathcal{I}_{K}(\mathbf{h}_{v}^{L})\right]$
@@
$p_{i}(G)=1\text{ if }\exists v\in V_{G},\,p_{i}(v)=1$
$p_{i}(G)=0\text{ if }\forall v\in V_{G},\,p_{i}(v)=0$
To better illustrate the process of identifying hidden predicates, we present a simple example in Figure LABEL:fig:graph_a. This scenario involves a binary graph classification task, a common setup in GNN applications. In this example, we have five input graphs, with each node characterized by two attributes: degree and type. The types, “A” and “B”, are encoded as 0 and 1, respectively. A GNN with a single message-passing layer is applied, generating a 2-dimensional embedding for each node (i.e., $d_{L}=2$). As only one message-passing layer is used, structural patterns are extracted based on the nodes and their first-order neighbors.
@@ 3.2 Determining the Logical Structure of $\phi_{M}$
In this subsection, we aim to construct global logical rules $\phi_{M}$ based on hidden predicates $P$ for each class, which serve the same functionality as the original GNN $M$. To achieve this, we adopt a data-driven approach. We process all training instances from class $c\in C$ that are correctly predicted by $M$, evaluating them against the predicates $P$ and recording their respective activation patterns. The results are stored in a binary matrix $\Phi_{c}$ for each class $c$, where the columns correspond to the predicates in $P$, and the rows represent the training instances correctly classified as $c$ by $M$. Specifically, an entry $\Phi_{c}[i,j]=1$ denotes that the $j$-th instance exhibits the $i$-th predicate, while $\Phi_{c}[i,j]=0$ indicates otherwise, as illustrated in Figure LABEL:fig:graph_b.
@@
$\forall G,\text{ }$ $\left(p_{1}\land p_{2}\land p_{4}\right)\lor\left(p_{1}\land p_{5}\right)\Rightarrow\textit{label}(G,0),$
$\forall G,\text{ }$ $\left(p_{1}\land p_{2}\land p_{3}\right)\lor\left(p_{1}\land p_{3}\land p_{4}\land p_{6}\right)\Rightarrow\textit{label}(G,1).$
Here, we omit $G$ in $p_{j}(G)$ when the context is clear. In addition to the descriptive rules $\bar{\phi}_{M}^{c}$, we also record and learn the recurring connectivity patterns of predicates for each conjunctive form. For instance, in the case of $(p_{1}\land p_{2}\land p_{4})$, both $p_{1}$ and $p_{2}$ are connected to $p_{4}$, which can be represented using an adjacency matrix. Multiple connectivity patterns may exist, and these can be learned in a data-driven manner, similar to how $\bar{\phi}_{M}$ is learned. We denote the collection of connectivity patterns as $\psi_{M}$, which is useful for graph generation and motif-level rule grounding.
@@
$\forall G,\ \neg p_{3}(G)\Rightarrow\textit{label}(G,0),\quad\forall G,\ p_{3}(G)\Rightarrow\textit{label}(G,1).$
@@ 3.3 Grounding $\phi_{M}$ into the Input Space
The next challenge is interpreting $\phi_{M}$. To address this, we focus on grounding its building blocks—predicates $P$—into the input space $\mathbf{X}$, thereby bridging the abstract logic with tangible input features. Recall that a predicate $p_{j}$ encodes a subgraph centered on a node $v$, encompassing its $1,\dots,L$-hop neighborhoods. Each neighbour node is also mapped to some predicate by the predicate function $f$. Thus, we can leverage the input features of the node $v$ along with its neighbour predicates to learn rules that relate the predicate $p_{j}$ to the input space. To achieve this, we define the input feature for the subgraph centered at $v$ as $\mathbf{Z}_{v,L}$, which is constructed by concatenating the information of nodes at each neighborhood level $l$:
$\mathbf{Z}_{v,L}=\text{CONCAT}\left(\mathbf{X}_{v},\text{Encode}\left(\{f(u)\mid u\in\mathcal{N}^{(1)}(v)\}\right),\right.$
$\left.\dots,\text{Encode}\left(\{f(u)\mid u\in\mathcal{N}^{(L)}(v)\}\right)\right)$
where $\mathcal{N}^{(l)}(v)$ represents the set of $l$-hop neighbors of $v$, and Encode transforms the collection of predicates into a frequency-based encoding, serving as an order-invariant aggregation operator applied to the predicates of the $l$-hop neighbors. For example, as illustrated in Figure LABEL:fig:graph_c, the subgraph input for node 1 is $\mathbf{Z}_{1,1}=(1,A,p_{4})$, which indicates that node 1 has a degree of 1 and a type of “A”. It also connects to predicate $p_{4}$, representing its sole neighbor, node 2. Similarly, the subgraph input for node 2 is $\mathbf{Z}_{2,1}=(3,B,p_{1},p_{1},p_{2})$. Here, we omit the encoding in $\mathbf{Z}$ for demonstration purposes.
Since model-agnosticism is a key requirement for LogicXGNN, we approximate the message-passing layers using interpretable models. In this work, we utilize decision trees for this purpose. Specifically, decision trees are employed to generate input-level rules that distinguish predicates with isomorphic subgraphs but different embedding patterns. For this task, the training data for each predicate label $j$ is the set of subgraph inputs of the nodes $v$ that activate $p_{j}$, represented as $\{\mathbf{Z}_{v,L}\mid p_{j}(v)=1\}$. For example, the training data for $p_{1}$ (identified as (“cde85e”, 0)) is $\{\mathbf{Z}_{1,1},\mathbf{Z}_{3,1},\dots,\mathbf{Z}_{22,1}\}$, while the training data for $p_{2}$ (identified as (“cde85e”, 1)) is $\{\mathbf{Z}_{4,1},\mathbf{Z}_{11,1},\mathbf{Z}_{20,1}\}$. The decision tree generates input-level rules such as $z[1]\leq 0.5$ for $p_{1}$, and otherwise for $p_{2}$. Given the prior knowledge that $z[1]$ represents the node’s type, we recognize that $p_{1}$ encodes whether the node is of type “A”, and $p_{2}$ is true when the node is of type “B”.
@@
$\forall G,\,\forall v\in V_{G}\,\left(\text{degree}(v)\neq 2\right)$ $\Rightarrow\textit{label}(G,0),$
$\forall G,\,\exists v\in V_{G}\,\left(\text{degree}(v)=2\right)$ $\Rightarrow\textit{label}(G,1).$
Sometimes, it is necessary to include neighbor predicates for effective grounding. For instance, $p_{4}$, supported by $\{\mathbf{Z}_{2,1},\mathbf{Z}_{12,1}\}$, encodes the presence of predicate $p_{2}$ as a neighbor. Since $p_{2}$ encodes type “B,” $p_{4}$ can be interpreted as having a neighbor node of type “B”. By incorporating connectivity patterns $\psi_{M}$, we can generate motif-level rules rather than node-level through the grounding of adjacent predicates. For example, in our simple GNN case, as shown in Figure LABEL:fig:graph_c, $\psi_{M}$ learns that two instances of $p_{3}$ are connected and share connections with the same other predicates. Given that these two nodes corresponding to $p_{3}$ have a degree of 2, we can directly infer the following rule:
$\forall G,\,\neg\text{has\_cycle}(G)\Rightarrow\textit{label}(G,0),$
$\forall G,\,\text{has\_cycle}(G)\Rightarrow\textit{label}(G,1)$

### v2

@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as powerful tools for modeling and analyzing graph-structured data, achieving remarkable performance across diverse domains, including drug discovery [28, 12, 21], fraud detection [19], and recommender systems [5]. Despite their success, GNNs share the black-box nature inherent to the neural network family, which poses challenges to their further development in high-reliability applications such as healthcare [1, 4].
@@
To this end, several explainability methods have been developed to uncover the inner decision-making mechanisms of GNNs. However, most of these methods are limited to providing local explanations tailored to specific input instances or rely on interpretations based on input feature attributions [18, 31, 24, 13, 22]. Another line of research focuses on global explanations that describe the overall behavior of models [30, 3, 2]. These approaches offer more human-readable and precise explanations by leveraging logical formulas and interpretable concepts. However, they have limitations, such as relying on predefined concepts and generating rules that explain only a limited set of patterns within each class without effectively distinguishing between these classes.
@@
To this end, we propose LogicXGNN, a novel framework that extracts interpretable logic rules to explain the internal reasoning process of GNNs while maintaining classification accuracy comparable to the original model. LogicXGNN ensures interpretability by grounding the hidden predicates of the rules into the input space using decision trees, which are both computationally efficient and data-driven. To the best of our knowledge, LogicXGNN is the first interpretable, rule-based functional equivalent of GNNs. Additionally, it can function as a generative model, with notable potential in fields such as drug design and knowledge discovery. We demonstrate the effectiveness of our approach through extensive experiments on real-world benchmarks, including the IMDB [16], Mutagenicity [6], and BBBP [27] datasets. In summary, we make the following contributions:
  • We propose LogicXGNN, a novel post-hoc framework for extracting interpretable logic rules from GNNs. It models the detailed computational processes in GNNs—such as message passing and pooling—making it model-agnostic, efficient, and fully data-driven.
  • Experimental results demonstrate that LogicXGNN significantly outperforms state-of-the-art global explanation methods in preserving the discriminative power of GNNs, while achieving 10 to 100 times faster overall runtime. As a rule-based classifier, LogicXGNN can even surpass the original GNNs on well-structured datasets, such as molecular graphs.
  • Thanks to its high interpretability and strong alignment with the decision-making processes of GNNs, LogicXGNN supports effective knowledge discovery—it can extract detailed and accurate knowledge (e.g., chemical) that is often overlooked by existing approaches.
@@
  • As LogicXGNN makes decision-making in GNNs transparent, it can be used as a generative model for creating graph instances, holding significant potential in fields such as drug design.
@@ 2.1 Graph Neural Networks
$\mathbf{h}_{v}^{l+1}=\text{UPD}\left(\mathbf{h}_{v}^{l},\text{AGG}\left(\left\{\text{MSG}(\mathbf{h}_{v}^{l},\mathbf{h}_{u}^{l})\mid\mathbf{A}_{uv}=1\right\}\right)\right),$ (1)
where $\mathbf{h}_{v}^{0}=\mathbf{X}_{v}$ is the initial feature vector of node $v$, and $\mathbf{h}_{v}^{l}$ represents the final node embedding at the final layer $L$. The update function UPD, aggregation operation AGG, and message function MSG define the architecture of a GNN. For instance, Graph Convolutional Networks (GCN) [10] use an identity message function, mean aggregation, and a weighted update. In essence, the GNN $\mathcal{M}$ aggregates information from both the feature space and the topological structure of $G$ to compute node embeddings, which are then optimized for downstream tasks such as graph classification.
@@
Suppose we have a set of graphs $\mathcal{G}$ and a label function $f:\mathcal{G}\rightarrow\{1,\dots,C\}$ that assigns one of $C$ classes to each graph in $\mathcal{G}$. To approximate $f$, we define a GNN model $\mathcal{M}$ for graph classification by passing the graph embeddings $\mathbf{h}_{G}^{L}$ to a fully connected layer followed by a softmax function. Here, the graph embeddings are commonly computed by taking the mean of all node embeddings in the graph $\mathbf{h}_{G}^{L}:=\text{mean}(\mathbf{h}_{v}^{L}\mid v\in V_{G})$ through the operation global_mean_pooling.
@@ 2.2 First-Order Logic Rules for GNN Interpretability
First-order logic (FOL), also known as predicate logic, is highly interpretable to humans, making it an excellent tool for explaining the behaviour of neural networks [34]. In this paper, our proposed framework, LogicXGNN, aims to elucidate the inner decision-making process of a GNN $\mathcal{M}$ using a Disjunctive Normal Form (DNF) formula $\phi_{M}$. The formula $\phi_{M}$ is a logical expression that can be described as a disjunction of conjunctions (OR of ANDs) over a set of predicates $P$, where each $p_{j}$ represents a logical condition or property defined on the graph structure $\mathbf{A}$ and input features $\mathbf{X}$. Importantly, $\phi_{M}$ incorporates the universal quantifier ($\forall$), providing a global explanation that is specific to a class of instances. This makes $\phi_{M}$ a rule-based model that is functionally equivalent to $\mathcal{M}$ a significant advantage not typically offered by other explanation work. To provide a concrete example of $\phi_{M}$, suppose the model $M$ is trained to determine whether a molecule $G$ is soluble. In this case, we may extract the following logical rules $\phi_{M}$ from $M$:
$\forall G,\,(p_{1}(G)\land p_{2}(G))\lor(p_{3}(G)\land p_{4}(G))\Rightarrow\text{label}(G,s),$ (2)
where $p_{1}(G)$ represents “The molecule contains a hydroxyl group (-OH)”, $p_{2}(G)$ represents “The molecule has a ring structure”, $p_{3}(G)$ represents “The molecule contains a carbonyl group (C=O)”, $p_{4}(G)$ represents “The molecule has a high degree of branching”, and $\text{label}(G,s)$ is a predicate indicating that graph $G$ is assigned the class label $s$ (“soluble”).
@@ 3 The LogicXGNN Framework
(a) Identifying hidden predicates $P$ for $\phi_{M}$.
(b) Extracting $\phi_{M}$.
(c) Grounding $\phi_{M}$.
Figure 1: An overview of the LogicXGNN framework, which involves identifying hidden predicates, extracting rules, and grounding these rules in the input space for interpretability.
In this section, we show how LogicXGNN (denoted as $\phi_{M}$) addresses the aforementioned challenges. For ease of discussion, we divide LogicXGNN into three critical sub-problems. Figure 1 illustrates an overview of the LogicXGNN framework.
@@ 3.1 Identifying Hidden Predicates $P$ for $\phi_{M}$
We start by discussing the identification of hidden predicates for graph classification tasks. In fact, since we define hidden predicates at the node level, our framework naturally extends to node classification tasks, as demonstrated in Appendix A.9.
As previously mentioned, the desired predicates $P$ should capture commonly shared patterns in both graph structures $\mathbf{A}$ and hidden embeddings $\mathbf{h}^{L}$ across a set of instances in the context of GNNs. While graph structure information can be encoded into hidden embeddings, it often becomes indistinguishable due to oversmoothing during the message-passing process [11, 29]. To mitigate this, we explicitly model common patterns in graph structures.
@@
$\text{Pattern}_{\text{struct}}(v)=\text{Hash}\big{(}\text{ReceptiveField}(v,\mathbf{A},L)\big{)}.$ (3)
Next, we discuss common patterns in the hidden embeddings. During the training of GNNs for classification tasks, the hidden embeddings are optimized to differentiate between classes. Empirically, we find that a small subset of specific dimensions in the final-layer embeddings $\mathbf{h}_{G}^{L}$ is sufficient to distinguish instances from different classes when using appropriate thresholds, often achieving similar accuracy to the original neural networks. Similar observations have been reported in [7].
@@
$\text{DecisionTree}(\{\mathbf{h}_{G}^{L}\mid G\in\mathcal{G}\},\mathbf{Y})\rightarrow(K,T)$ (4)
@@
$\mathcal{I}_{k}(\mathbf{h}_{v}^{L})=1\text{ if }\mathbf{h}_{v}^{L}[k]\geq T_{k},\text{ else }0$ (5)
@@
$\text{Pattern}_{\text{emb}}(v)=\left[\mathcal{I}_{1}(\mathbf{h}_{v}^{L}),\mathcal{I}_{2}(\mathbf{h}_{v}^{L}),\dots,\mathcal{I}_{K}(\mathbf{h}_{v}^{L})\right]$ (6)
@@
$p_{i}(G)=1\text{ if }\exists v\in V_{G},\,p_{i}(v)=1,\quad p_{i}(G)=0\text{ if}\forall v\in V_{G},\,p_{i}(v)=0.$ (7)
To better illustrate the process of identifying hidden predicates, we present a simple example in Figure 1(a). This scenario involves a binary graph classification task, a common setup in GNN applications. In this example, we have five input graphs, with each node characterized by two attributes: degree and type. The types, “A” and “B”, are encoded as 0 and 1, respectively. A GNN with a single message-passing layer is applied, generating a 2-dimensional embedding for each node (i.e., $d_{L}=2$). As only one message-passing layer is used, structural patterns are extracted based on the nodes and their first-order neighbors.
@@ 3.2 Determining the Logical Structure of $\phi_{M}$
In this subsection, we aim to construct global logical rules $\phi_{M}$ based on hidden predicates $P$ for each class, which serve the same functionality as the original GNN $M$. To achieve this, we adopt a data-driven approach. We process all training instances from class $c\in C$ that are correctly predicted by $M$, evaluating them against the predicates $P$ and recording their respective activation patterns. The results are stored in a binary matrix $\Phi_{c}$ for each class $c$, where the columns correspond to the predicates in $P$, and the rows represent the training instances correctly classified as $c$ by $M$. Specifically, an entry $\Phi_{c}[i,j]=1$ denotes that the $j$-th instance exhibits the $i$-th predicate, while $\Phi_{c}[i,j]=0$ indicates otherwise, as illustrated in Figure 1(b).
@@
$\forall G,\left(p_{1}\land p_{2}\land p_{4}\right)\lor\left(p_{1}\land p_{5}\right)\Rightarrow\textit{label}(G,0),\quad\forall G,\left(p_{1}\land p_{2}\land p_{3}\right)\lor\left(p_{1}\land p_{3}\land p_{4}\land p_{6}\right)\Rightarrow\textit{label}(G,1).$ (8)
Here, we omit $G$ in $p_{j}(G)$ when the context is clear. In addition to the descriptive rules $\bar{\phi}_{M}^{c}$, we also record the recurring connectivity patterns of predicates for each conjunctive form. For instance, in the case of $(p_{1}\land p_{2}\land p_{4})$, both $p_{1}$ and $p_{2}$ are connected to $p_{4}$, which can be represented using an adjacency matrix. Multiple connectivity patterns may exist, and these can be learned in a data-driven manner, similar to how $\bar{\phi}_{M}$ is learned. We denote the collection of connectivity patterns as $\psi_{M}$, which is useful for graph generation and motif-level rule grounding. Additional details on graph generation are provided in Appendix A.11.
@@
$\forall G,\ \neg p_{3}(G)\Rightarrow\textit{label}(G,0),\quad\forall G,\ p_{3}(G)\Rightarrow\textit{label}(G,1).$ (9)
@@ 3.3 Grounding $\phi_{M}$ into the Input Space
The next challenge is interpreting $\phi_{M}$. To address this, we focus on grounding its building blocks—predicates $P$—into the input space $\mathbf{X}$, thereby bridging the abstract logic with tangible input features. Recall that a predicate $p_{j}$ encodes a subgraph centered on a node $v$, encompassing its $1,\dots,L$-hop neighborhoods. Each neighbour node is also mapped to some predicate by the predicate function $f$. Thus, we can leverage the input features of the node $v$ along with its neighbour predicates to infer rules that relate the predicate $p_{j}$ to the input space. To achieve this, we define the input feature for the subgraph centered at $v$ as $\mathbf{Z}_{v,L}$, which is constructed by concatenating the information of nodes at each neighborhood level $l$:
$\mathbf{Z}_{v,L}=\text{CONCAT}\left(\mathbf{X}_{v},\text{ENCODE}\left(\{f(u)\mid u\in\mathcal{N}^{(1)}(v)\}\right),\right.\left.\dots,\text{ENCODE}\left(\{f(u)\mid u\in\mathcal{N}^{(L)}(v)\}\right)\right)$ (10)
where $\mathcal{N}^{(l)}(v)$ represents the set of $l$-hop neighbors of $v$, and ENCODE transforms the collection of predicates into a frequency-based encoding, serving as an order-invariant aggregation operator applied to the predicates of the $l$-hop neighbors. For example, as illustrated in Figure 1(c), the subgraph input for node 1 is $\mathbf{Z}_{1,1}=(1,A,p_{4})$, which indicates that node 1 has a degree of 1 and a type of “A". It also connects to predicate $p_{4}$, representing its sole neighbor, node 2. Similarly, the subgraph input for node 2 is $\mathbf{Z}_{2,1}=(3,B,p_{1},p_{1},p_{2})$. Here, we omit the encoding in $\mathbf{Z}$ for demonstration purposes.
Since model-agnosticism is a key requirement for LogicXGNN, we approximate the message-passing layers using interpretable models. In this work, we utilize decision trees for this purpose. Specifically, decision trees are employed to generate input-level rules that distinguish predicates with isomorphic subgraphs but different embedding patterns. For this task, the training data for each predicate label $j$ is the set of subgraph inputs of the nodes $v$ that activate $p_{j}$, represented as $\{\mathbf{Z}_{v,L}\mid p_{j}(v)=1\}$. For example, the training data for $p_{1}$ (identified as (“cde85e”, 0)) is $\{\mathbf{Z}_{1,1},\mathbf{Z}_{3,1},\dots,\mathbf{Z}_{22,1}\}$, while the training data for $p_{2}$ (identified as (“cde85e”, 1)) is $\{\mathbf{Z}_{4,1},\mathbf{Z}_{11,1},\mathbf{Z}_{20,1}\}$. The decision tree then generates input-level rules such as $\mathbf{Z}[1]\leq 0.5$ for $p_{1}$, and the opposite condition for $p_{2}$. Recall that $\mathbf{Z}[1]$, i.e., the first dimension of the feature vector $\mathbf{Z}$, encodes the node type. Given this, we recognize that $p_{1}$ indicates that the node is of type “A”, while $p_{2}$ indicates that the node is of type “B”.
@@
$\forall G,\,\forall v\in V_{G}\,\left(\text{degree}(v)\neq 2\right)\Rightarrow\textit{label}(G,0),\quad\forall G,\,\exists v\in V_{G}\,\left(\text{degree}(v)=2\right)\Rightarrow\textit{label}(G,1).$ (11)
Sometimes, it is necessary to include neighbor predicates for effective grounding. For instance, $p_{4}$, supported by $\{\mathbf{Z}_{2,1},\mathbf{Z}_{12,1}\}$, encodes the presence of predicate $p_{2}$ as a neighbor. Since $p_{2}$ encodes type “B”, $p_{4}$ can be interpreted as having a neighbor node of type “B”. By incorporating connectivity patterns $\psi_{M}$, we can generate motif-level rules rather than node-level through the grounding of adjacent predicates. For example, in our simple GNN case, as shown in Figure 1(c), $\psi_{M}$ learns that two instances of $p_{3}$ are connected and share connections with the same other predicates. Given that these two nodes corresponding to $p_{3}$ have a degree of 2, we can directly infer the following rule:
$\forall G,\,\neg\text{has\_cycle}(G)\Rightarrow\textit{label}(G,0),\quad\forall G,\,\text{has\_cycle}(G)\Rightarrow\textit{label}(G,1),$ (12)